OpenStudy (anonymous):

Find the vertical asymptotes, if any, of the graph of the rational function. i will post the function and answer choices

OpenStudy (anonymous):

\[x-4 \over x(x-4) \] answer choices A. x = 4 and x = 4 B. x = 4 C. x = 0 and x = 4 D. no vertical asymptote

OpenStudy (anonymous):

in order to find vertical Find asymptotes find at which value the denominator get zero in this case x(x-4)=0 so x=0 x-4=0 x=4 x=0 and x=4 option C

OpenStudy (anonymous):

thnx